OpenStudy (anonymous):

For a triangle with side lengths 1.5 feet, 2 feet and 3 feet, the area of the triangle is:

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Are you Here Radar?

OpenStudy (radar):

Use the formula for calculating area of a triangle when the three sides are known.. First compute s. s is the value of the perimeter divided by 2. In this problem the sum of the sides are: 1.5+2+3=6.5 ft. So s=3.25

OpenStudy (anonymous):

Thanks

OpenStudy (radar):

Yes I am here, are we having fun yet?

OpenStudy (anonymous):

No, geometry is hard, I have this question on a quiz, it does not list this answer.

OpenStudy (radar):

Here is the formula:\[A=\sqrt{s(s-a)(s-b)(s-c)}\] Where a, b, c are the sides

OpenStudy (anonymous):

Are you sure, that is the answer?

OpenStudy (radar):

That is the formula. plug in values s=3.25, a=1.5, b=2, and c=3.

OpenStudy (anonymous):

ok, maybe it is none of them, since I do not show your answer.

OpenStudy (radar):

\[\sqrt{3.25(3.25-1.5)(3.25-2)(3.25-3)}\]

OpenStudy (anonymous):

I am having a hard time with this open book test lol, they are the worst

OpenStudy (radar):

\[\sqrt{3.25(1.75)(1.25)(0.25)}\]

OpenStudy (radar):

using my calculator i get:\[\sqrt{1.77734375}\]

OpenStudy (anonymous):

ok

OpenStudy (radar):

that would be: 1.33317 sq ft

OpenStudy (radar):

Better double check my figures cdunn

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I think you are right

OpenStudy (anonymous):

I have 19 replies, how can I see the rest of them?

OpenStudy (radar):

are you looking at the open study web page, I am and I see all 20 replies

OpenStudy (anonymous):

I cannot see them

OpenStudy (radar):

maybe if you refresh your browser

OpenStudy (anonymous):

strange, I cannot see them, but their answer is the same as your answer?

OpenStudy (radar):

ok good luck with it. I got that formula for area from a very old text. bye

OpenStudy (anonymous):

thanks